OpenStudy (anonymous):

Csc^2 theta = 3Sec^2 theta. How do convert either Csc to sec or Sec to csc? Also how do you solve it in radians?

OpenStudy (anonymous):

i would start with \[\sin^2(\theta)=\frac{\cos^2(\theta)}{3}\] by taking reciprocals

OpenStudy (anonymous):

or maybe go right to \[\tan^2(\theta)=\frac{1}{3}\]

OpenStudy (anonymous):

Sorry I don't really understand how you got to that response :/ I'm new to this.

OpenStudy (anonymous):

ok i can't work with cosecant , secant etc lets go step by step and work with since and cosine

OpenStudy (anonymous):

since \(\csc(x)=\frac{1}{\sin(x)}\) and \(\sec(x)=\frac{1}{\cos(x)}\) your original equation is \[\frac{1}{\sin^2(x)}=\frac{3}{\cos^2(x)}\]

OpenStudy (anonymous):

solve as any ratio first, get \[\cos^2(x)=3\sin^2(x)\]

OpenStudy (anonymous):

so far so good? only algebra at these steps no trig

OpenStudy (anonymous):

yeah, i didn't thing about changing back to sin and cos

OpenStudy (anonymous):

eventually you should come to \[\tan^2(x)=\frac{1}{3}\] making \[\tan(x)=\pm\frac{1}{\sqrt3}\]

OpenStudy (anonymous):

cos/3sin right?

OpenStudy (anonymous):

gives tan?

OpenStudy (freckles):

I would be careful about dividing both sides by 0 \[\cos^2(x)=3\sin^2(x) \\ 1-\sin^2(x)=3\sin^2(x) \\ 1=4\sin^2(x) \]

OpenStudy (freckles):

but given the initial problem sin(x) cannot be 0 and neither can cosine so you are good

OpenStudy (anonymous):

Yeah I know that cos^2 = 1-sin^2 but I didn't know you couldnt go it with it csc and sec

OpenStudy (freckles):

sat's way will work i'm just saying we need to be careful when we divide by 0 but really we aren't dividing by 0 because we had csc and sec in the initial problem

OpenStudy (anonymous):

Oh okay, I got you, Thanks :)

OpenStudy (anonymous):

I have found my answer and really appreciate your guy's help! It really means a lot to me :)

OpenStudy (freckles):

congrats